# From Image Denoisers to Regularizing Imaging Inverse Problems: An Overview

### v1

@@ 2.1 Foundations of Proximal Calculus
In the case where $f$ is differentiable, the proximal operator may be interpreted as a backwards Euler discretization of gradient flow $\dot{x}(t)=-\nabla f(x(t))$. With step size $\eta=\lambda$, this is given by $x_{k+1}=x_{k}-\eta\nabla f(x_{k+1})$, which is precisely the first-order optimality condition for 6. The proximal operator generalizes this concept to nonsmooth and $\infty$-valued functions. Furthermore, the proximal operator is a generalization of projection onto a convex set. Suppose $f=\chi_{C}$ is the characteristic function of a convex set $C$, taking values 0 in $C$ and $\infty$ otherwise. Then, the proximal operator $\operatorname{prox}_{\lambda f}$ is precisely the (Euclidean) projection onto $C$. This equivalence is useful when interpreting constrained optimization.
@@ 2.1.1 Monotone Operators
For a (proper and closed) convex function $f$, the subdifferential operator $\partial f$ can be seen to be a monotone operator (and indeed, something stronger called cyclically monotone) rockafellar1966characterization . Analogously to the backward Euler interpretation above, the proximal map can be equivalently characterized as the resolvent of the subdifferential operator (rockafellar2009variational, , Sec. 12.C.):
$\operatorname{prox}_{\lambda f}=J_{\partial f}\coloneqq(I+\lambda\,\partial f)^{-1}.$ (7)
The resolvent is monotone for proper closed functions, and maximally monotone if $f$ is additionally convex. This identity reveals that proximal operations are fundamentally linked to the theory of maximal monotone operators. These theoretical foundations underpin the development of proximal algorithms, particularly the proximal gradient method, arising in convex optimization by leveraging results from monotone operator theory.
@@ 2.1.2 Composite Optimization and Operator Splitting
For a Hilbert space $\mathcal{X}$, let $A:\mathcal{X}\rightarrow\mathcal{X}$ be $\beta$-cocoercive (i.e. $\langle x-y,Ax-Ay\rangle\geq\beta\|Bx-By\|^{2}$) for some $\beta>0$, and $B:\mathcal{X}\rightrightarrows\mathcal{X}$ be maximally monotone. Let $\lambda\in(0,2\beta)$ and set $\delta=\min\{1,\,\beta/\lambda\}+1/2$. Suppose that $\operatorname*{zer}(A+B)\neq\emptyset$, and define the forward-backward iterations as
$\begin{cases}y_{n}=x_{n}-\lambda Ax_{n},\\ x_{n+1}=J_{\lambda B}y_{n}-x_{n}.\end{cases}$ (8)
@@
    $(x_{n})_{n\in\mathbb{N}}$ converges weakly to a point in $\operatorname*{zer}(A+B)$;
@@
For a Hilbert space $\mathcal{X}$, let $A,B:\mathcal{X}\rightrightarrows\mathcal{X}$ be maximally monotone operators such that $\operatorname*{zer}(A+B)\neq\emptyset$. Let $\lambda>0$ be a step size and $x_{0}\in\mathbb{R}^{n}$ be an initialization. Consider the iterations
@@
$x^{k+1}=J_{\lambda B}(2J_{\lambda A}-I)x^{k}+(I-J_{\lambda A})x^{k}.$ (10)
@@
    $y-z_{n}$ converges strongly to zero,
@@
    $y$ and $z_{n}$ converge weakly to $J_{\lambda A}(x)$.
Note that in the case where the Hilbert space $\mathcal{X}$ is finite-dimensional, weak convergence is equivalent to strong convergence. To obtain the corresponding convex optimization, simply let $A$ and $B$ be the subdifferentials of some proper closed convex functions $f$ and $g$. We get convergence to a zero of $J_{A+B}$, equivalently a fixed point of $\operatorname{prox}_{f+g}$, using only proximal operators or subgradients of $f$ and $g$ separately. Furthermore, the fixed point is a minimum of $f+g$.
By casting the above monotone inclusion problem in the scope of convex functions, with $A$ being a derivative and $B$ being a proximal operator, we can obtain splitting schemes that optimize the sum of two convex functions, where one of the functions is smooth.
@@ 2.1.3 PnP Proximal Gradient Descent
For casting 8 into the context of convex optimization by letting $A=\nabla f$, $B=\partial g$, we consider the composite optimization problem:
@@
Let $f$ be $\mu$-strongly convex and $L$-smooth for some $\mu\geq 0,L>0$, and let $g$ be convex. For step size $\lambda\in(0,2/L)$, the PGD iterates satisfy
$\|x_{k}-x^{*}\|_{2}\leq\varrho^{k}\|x^{0}-x^{*}\|_{2},$ (13)
where $\varrho=\max(|1-\lambda L|,|1-\lambda\mu|)<1$. When $f$ is merely convex, the objective error decays as $\mathcal{O}(1/k)$. Moreover, the minimum of the residuals satisfies
@@ 2.1.4 Relaxed PnP Proximal Gradient Descent
The PGD iterations can be relaxed to accommodate for weakly convex functions $g$ while minimizing $f+g$. The relaxed iterations, for $\alpha\in(0,1)$, are hurault2024convergent ; chambolle2016ergodic :
@@
Let $f$ be convex and $L$-smooth, and $g$ be $M$-weakly convex. Then for $\alpha\in(0,1)$ and $\lambda<\min((\alpha L_{f})^{-1},\alpha M^{-1})$, define the objective $F=\lambda f+g$. The iterates $x_{k}$ in 15 satisfy
@@
    $F(x_{k})+\frac{\alpha}{2\lambda}\left(1-\frac{1}{\alpha}\right)^{2}\|x_{k}-x_{k+1}\|^{2}$ is non-increasing and convergent;
@@
    The sequence $(x_{k})$ has finite length, and $\min_{l<K}\|x_{l+1}-x_{l}\|=\mathcal{O}(1/\sqrt{k})$;
@@ 2.1.5 PnP Douglas–Rachford Splitting
For proper, convex, and closed functions $f$ and $g$, where $f$ is differentiable, one can substitute $A=\nabla f,B=\partial g$ into 9 to yield the update
$x_{k+1}=\operatorname{prox}_{\lambda g}(2\operatorname{prox}_{\lambda f}-I)x_{k}+(I-\operatorname{prox}_{\lambda f})x_{k}.$ (16)
@@
The residuals in the DRS iteration, given by
@@
The asymmetry of the Douglas–Rachford splitting gives rise to two possible splittings by switching the roles of $f$ and $g$ hurault2022proximal . In split form, it gives rise to PnP-DRS and PnP-DRSdiff, defined as follows, where the latter requires that $f$ is differentiable for convergence. Notably, the two variants of PnP with the DRS have slightly different convergence assumptions on $g$. The PnP-DRS iterations are given by
@@ 2.2 ADMM: Constrained Optimization to Operator Splitting
While PGD and DRS consider composite optimization in one variable, a more general problem may include equality constraints, such as in a Lagrangian. Alternatively, one may be interested in a problem of the form $f(x)+g(Kx)$ for some forward operator $K$ with a computable adjoint $K^{*}$. The alternating direction method of multipliers (ADMM) algorithm boyd_admm_book is equipped to address problems with a separable structure having the general form
$\min_{x,z}f(x)+g(z)\quad\text{subject to}\quad Kx+Bz=c.$ (17)
ADMM can be derived from applying DRS to a dual formulation gabay1983chapter ; ryu2016primer . To see this, introduce the augmented Lagrangian according to the equality constraint:
$\mathcal{L}_{\rho}(x,z,u)=f(x)+g(z)+u^{\top}(Kx+Bz-c)+\frac{\rho}{2}\|Kx+Bz-c\|_{2}^{2}.$ (18)
@@
$\begin{cases}x_{k+1}=\operatorname*{arg\,min}_{x}\,\mathcal{L}_{\rho}(x,z_{k},u_{k}),\\ z_{k+1}=\operatorname*{arg\,min}_{z}\,\mathcal{L}_{\rho}(x_{k+1},z,u_{k}),\\ u_{k+1}=u_{k}+\rho(Kx_{k+1}+Bz_{k+1}-c).\end{cases}$ (19)
The connection to Douglas–Rachford splitting emerges when considering the dual problem. We demonstrate this in the special case where B=IB=-I and $c=0$. In this case, the primal problem becomes
@@
$\max_{u}-f^{*}(-K^{*}u)-g^{*}(u),$ (21)
where $K^{*}$ is the adjoint of the operator $K$, equivalently, the matrix transpose in finite dimensions. The optimality conditions for the dual problem are:
$0$ $\in-A\partial f^{*}(-K^{*}u^{*})+\partial g^{*}(u^{*})$
@@
$\exists x^{*},z^{*}\ \text{s.t.}\quad z^{*}$ $=Kx^{*}\quad\text{where}\quad x^{*}\in\partial f^{*}(-A^{\top}u^{*}),\,z^{*}\in\partial g^{*}(u^{*}).$
@@
$T_{1}(u)$ $=-K\partial f^{*}(-K^{*}u)=\partial(f^{*}\circ(-K^{*}))(u),$
@@
We need to solve the inclusion problem $0\in T_{1}(u)+T_{2}(u)$. Using DRS, the iterations may be rewritten as
@@
The resolvent of $T_{2}$ may be recognized exactly as a proximal operator $J_{\rho T_{2}}=J_{\rho\partial g^{*}}=\operatorname{prox}_{\rho g^{*}}=I-\rho\operatorname{prox}_{\rho^{-1}g}(\rho^{-1}\cdot)$. The resolvent of the first operator may also be computed as
$J_{\rho T_{1}}(v)$ $=\operatorname{prox}_{\rho f^{*}\circ(-K^{*})}(v)$
$=v-\rho\operatorname{prox}_{\rho^{-1}[f^{*}\circ(-K^{*})]^{*}}(\rho^{-1}v)$
$=v-\rho\operatorname*{arg\,min}_{x}\inf_{z,-Az=x}f(z)+\frac{\rho}{2}\|x-\rho^{-1}v\|^{2}$
$=v+\rho K\operatorname*{arg\,min}_{z}\left(f(z)+\frac{\rho}{2}\|Az+\rho^{-1}v\|^{2}\right).$
@@
$\begin{cases}\hat{x}_{k+1,1}=\operatorname*{arg\,min}_{x_{1}}\left(f(x_{1})+z_{k}^{\top}(Kx_{1})+\frac{\rho}{2}\|Kx_{k,1}-\hat{x}_{k,2}\|^{2}\right),\\ \hat{x}_{k+1,2}=\operatorname*{arg\,min}_{x_{2}}\left(g(x_{2})-z_{k}^{\top}x_{2}+\frac{\rho}{2}\|Kx_{k+1,1}-\hat{x}_{k,2}\|^{2}\right),\\ z_{k+1}=z_{k}+\rho(K\hat{x}_{k+1,1}-\hat{x}_{k+1,2}).\end{cases}$ (22)
@@
We observe that the minimizations are precisely with respect to the first two arguments, exactly yielding the ADMM algorithm. In the context of imaging inverse problems, the PnP variant of the ADMM algorithm emerges by replacing the proximal operator (with respect to a nonsmooth regularizer $g$) with an off-the-shelf denoiser $D$, while minimizing the variational objective with an $\ell_{2}^{2}$ fidelity term: $\underset{x}{\min}\,\frac{1}{2}\|y-Kx\|_{2}^{2}+g(x)$. To solve this problem using ADMM, one applies the variable separation trick to reformulate the problem as
@@
$\begin{cases}x_{k+1}=\left(K^{*}K+\rho\,I\right)^{-1}\left(K^{*}y+\rho\,(z_{k}-u_{k})\right),\\ z_{k+1}=D\left(x_{k+1}+u_{k}\right),\\ u_{k+1}=u_{k}+\left(x_{k+1}-z_{k+1}\right).\end{cases}$ (PnP-ADMM)

### v2

@@ 2.1 Foundations of Proximal Calculus
In the case where $f$ is differentiable, the proximal operator may be interpreted as a backwards Euler discretization of gradient flow $\dot{x}(t)=-\nabla f(x(t))$. With step size $\eta=\lambda$, the proximal scheme $x_{k+1}=\operatorname{prox}_{\lambda f}(x_{k})$ is equivalently given by $x_{k+1}=x_{k}-\eta\nabla f(x_{k+1})$, which may be shown using the first-order optimality condition in 6. The proximal operator generalizes this concept to nonsmooth and $\infty$-valued functions. Furthermore, the proximal operator is a generalization of projection onto a convex set. Suppose $f=\chi_{C}$ is the characteristic function of a convex set $C$, taking values 0 in $C$ and $\infty$ otherwise. Then, the proximal operator $\operatorname{prox}_{\lambda f}$ is precisely the (Euclidean) projection onto $C$. This equivalence is useful when interpreting constrained optimization.
@@ 2.1.1 Monotone Operators
For a (proper and closed) convex function $f$, the subdifferential operator $\partial f$ can be seen to be a monotone operator (and indeed, something stronger called maximally cyclically monotone) rockafellar1966characterization . Analogously to the backward Euler interpretation above, the proximal map can be equivalently characterized as the resolvent of the subdifferential operator (rockafellar2009variational, , Sec. 12.C.):
$\operatorname{prox}_{\lambda f}=J_{\lambda\partial f}\coloneqq(\operatorname{Id}+\lambda\,\partial f)^{-1}.$ (7)
By Minty’s theorem, we have that the resolvent of a monotone operator is defined everywhere if and only if the monotone operator is maximally monotone (bauschke2011convex, , Thm. 21.1). Moreover, a function $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is firmly non-expansive, i.e. $\|T(x-y)\|^{2}+\|(\operatorname{Id}-T)(x-y)\|^{2}\leq\|x-y\|^{2}$, if and only if it is the resolvent of a maximally monotone operator (bauschke2011convex, , Cor. 23.8). This firm non-expansiveness condition is sufficient for the fixed point iteration $x_{n+1}=Tx_{n}$ to converge.
Other interesting formulas can be reformulated in terms of maximal monotone operators. For example, Moreau’s identity $\operatorname{prox}_{f}+\operatorname{prox}_{f^{*}}=\operatorname{Id}$ can be reformulated using resolvents as $\operatorname{Id}=J_{\gamma A}+\gamma^{-1}J_{\gamma^{-1}A^{-1}}\circ\gamma^{-1}\operatorname{Id}$ (bauschke2011convex, , Prop. 23.18). Another lesser known identity states: if $f$ is proper closed and convex, and $^{\gamma}\!f\coloneqq f\circ\operatorname{prox}_{f}$ is its Moreau envelope, then $\operatorname{prox}_{{}^{\gamma}\!f}(x)=x+(\gamma+1)^{-1}(\operatorname{prox}_{(\gamma+1)f}(x)-x)$ (bauschke2011convex, , Prop. 23.29). These theoretical foundations and relations underpin the development of proximal algorithms, particularly the proximal gradient method, arising in convex optimization by leveraging results from monotone operator theory.
@@ 2.1.2 Composite Optimization and Operator Splitting
In the following results, we denote the fixed points of an operator $A$ by $\operatorname{Fix}A\coloneqq\{x\in\mathcal{X}\mid Ax=x\}$, and the zeros of a (possibly set-valued) mapping $B$ by $\operatorname*{zer}B\coloneqq\{x\in\mathcal{X}\mid 0\in Bx\}$.
@@
For a Hilbert space $\mathcal{X}$, let $A:\mathcal{X}\rightarrow\mathcal{X}$ be $\beta$-cocoercive (i.e. $\langle x-y,Ax-Ay\rangle\geq\beta\|Ax-Ay\|^{2}$) for some $\beta>0$, and $B:\mathcal{X}\rightrightarrows\mathcal{X}$ be maximally monotone. Let $\lambda\in(0,2\beta)$ and $x_{0}\in\mathbb{R}^{n}$ be an initialization, and further set $\delta=\min\{1,\,\beta/\lambda\}+1/2$. Suppose that $\operatorname*{zer}(A+B)\neq\emptyset$, and define the forward-backward iterations as follows,
$\begin{cases}y_{k}=x_{k}-\lambda Ax_{k},\\ x_{k+1}=J_{\lambda B}y_{k}.\end{cases}$ (8)
@@
    $(x_{k})_{k\in\mathbb{N}}$ converges weakly to a point in $\operatorname*{zer}(A+B)$;
@@
For a Hilbert space $\mathcal{X}$, let $A,B:\mathcal{X}\rightrightarrows\mathcal{X}$ be maximally monotone operators such that $\operatorname*{zer}(A+B)\neq\emptyset$. Let $\lambda>0$ be a step size and $x_{0}\in\mathcal{X}$ be an initialization. Consider the iterations
@@
$x_{k+1}=J_{\lambda B}(2J_{\lambda A}-\operatorname{Id})x_{n}+(\operatorname{Id}-J_{\lambda A})x_{k}.$ (10)
@@
    $y-z_{k}$ converges strongly to zero,
@@
    $y$ and $z_{k}$ converge weakly to $J_{\lambda A}(x)$.
Note that in the case where the Hilbert space $\mathcal{X}$ is finite-dimensional, weak convergence is equivalent to strong convergence. The reflected resolvent of $A$ appears in 10; since $A$ is maximally monotone, the reflected resolvent is a non-expansive operator (bauschke2011convex, , Cor. 23.10), allowing for a contraction-like argument.
To obtain the corresponding optimization method, simply let $A$ and $B$ be the subdifferentials of some proper closed convex functions $f$ and $g$. We get convergence to a zero of A+B{A+B}, equivalently a fixed point of $\operatorname{prox}_{f+g}$, using only proximal operators or subgradients of $f$ and $g$ separately. Furthermore, the fixed point is a minimum of $f+g$.
FBS and DRS impose different requirements for efficient computation, where the former needs that $\nabla f$ and $\operatorname{prox}_{g}$ are easy to compute (and $\nabla f$ is Lipschitz, hence cocoercive by Baillon–Haddad), and the latter requires that both $\operatorname{prox}_{f}$ and $\operatorname{prox}_{g}$ are easy to compute. Moreover, DRS can be extended to finding a zero of a finite sum of maximally monotone operators, with the resulting algorithm known as the parallel splitting algorithm (bauschke2011convex, , Prop. 25.7). Sharp convergence rates for FBS and DRS applied to composite optimization can be found in davis2015convergence ; molinari2019convergence .
@@ 2.1.3 PnP Proximal Gradient Descent
By casting the above monotone inclusion problem 8 in the scope of convex functions, with $A$ being a derivative and $B$ being a proximal operator, we can obtain splitting schemes that optimize the sum of two convex functions, where one of the functions is smooth. Letting $A=\nabla f$, $B=\partial g$, we consider solving the following composite optimization problem:
@@
Let $f$ be $\mu$-strongly convex and $L$-smooth for some $\mu\geq 0,L>0$, and let $g$ be convex. For step size $\lambda\in(0,2/L)$ and initialization $x_{0}$, the PGD iterates satisfy
$\|x_{k}-x^{*}\|_{2}\leq\varrho^{k}\|x_{0}-x^{*}\|_{2},$ (13)
where $\varrho=\max(|1-\lambda L|,|1-\lambda\mu|)<1$, and $x^{*}$ is the minimizer of $f+g$. When $f$ is merely convex, the objective error decays as $\mathcal{O}(1/k)$. Moreover, the minimum of the residuals satisfies
@@ 2.1.4 Relaxed PnP Proximal Gradient Descent
The PGD iterations can be relaxed to accommodate for weakly convex functions $g$ while minimizing $f+g$. The relaxed iterations, for some relaxation parameter $\alpha\in(0,1)$, are hurault2024convergent ; chambolle2016ergodic :
@@
Let $f$ be convex and $L_{f}$-smooth, and $g$ be $M$-weakly convex. Then for $\alpha\in(0,1)$ and $\lambda<\min((\alpha L_{f})^{-1},\alpha M^{-1})$, define the objective $F=f+g$. The iterates $x_{k}$ in 15 satisfy
@@
    $F(x_{k})+\frac{\alpha}{2}\left(1-\frac{1}{\alpha}\right)^{2}\|x_{k}-x_{k-1}\|^{2}$ is non-increasing and convergent;
@@
    The sequence $(x_{k})$ has finite length, i.e., $\sum_{k\in\mathbb{N}}\|x_{k+1}-x_{k}\|<+\infty$. Moreover, $\min_{l<k}\|x_{l+1}-x_{l}\|=\mathcal{O}(1/\sqrt{k})$;
@@ 2.1.5 PnP Douglas–Rachford Splitting
For proper, convex, and closed functions $f$ and $g$, one can substitute $A=\partial f,B=\partial g$ into 9 to yield the update
$x_{k+1}=\operatorname{prox}_{\lambda g}(2\operatorname{prox}_{\lambda f}-\operatorname{Id})x_{k}+(\operatorname{Id}-\operatorname{prox}_{\lambda f})x_{k}.$ (16)
@@
The residuals in the DRS iteration 16, given by
@@
The asymmetry of the Douglas–Rachford splitting gives rise to two possible splittings by switching the roles of $f$ and $g$ hurault2022proximal . These two splittings give rise to PnP-DRS and PnP-DRSdiff, defined as follows. Notably, the two variants of PnP with the DRS have slightly different convergence assumptions on $g$, and with PnP-DRSdiff further requiring that $f$ is differentiable. The PnP-DRS iterations are given by
@@ 2.2 ADMM: Constrained Optimization to Operator Splitting
While PGD and DRS consider composite optimization in one variable, a more general problem may include equality constraints, such as in a Lagrangian. Alternatively, one may be interested in a problem of the form $f(x)+g(Kx)$ for some forward operator $K$ with a computable adjoint $K^{\top}$. The alternating direction method of multipliers (ADMM) algorithm boyd_admm_book is equipped to address problems with a separable structure having the more general form
$\min_{x,z}f(x)+g(z)\quad\text{subject to}\quad Kx+K^{\prime}z=c,$ (17)
where $K^{\prime}$ is another linear operator, and the slack variable $c$ may arise from e.g. data constraints arising from an underlying inverse problem. ADMM can be derived from applying DRS to a dual formulation gabay1983chapter ; ryu2016primer . To see this, introduce the augmented Lagrangian according to the equality constraint:
$\mathcal{L}_{\rho}(x,z,u)=f(x)+g(z)+u^{\top}(Kx+K^{\prime}z-c)+\frac{\rho}{2}\|Kx+K^{\prime}z-c\|_{2}^{2}.$ (18)
@@
$\begin{cases}x_{k+1}=\operatorname*{arg\,min}_{x}\,\mathcal{L}_{\rho}(x,z_{k},u_{k}),\\ z_{k+1}=\operatorname*{arg\,min}_{z}\,\mathcal{L}_{\rho}(x_{k+1},z,u_{k}),\\ u_{k+1}=u_{k}+\rho(Kx_{k+1}+K^{\prime}z_{k+1}-c).\end{cases}$ (19)
The connection to Douglas–Rachford splitting emerges when considering the dual problem. We demonstrate this in the special case where $K^{\prime}=-\operatorname{Id}$ and $c=0$. In this case, the primal problem becomes
@@
$\max_{u}-f^{*}(-K^{\top}u)-g^{*}(u),$ (21)
where $K^{\top}$ is the adjoint of the operator $K$, equivalently, the matrix transpose when $\mathcal{X}=\mathbb{R}^{n}$. The optimality conditions for the dual problem are:
$0$ $\in-K\partial f^{*}(-K^{\top}u^{*})+\partial g^{*}(u^{*})$
@@
$\exists x^{*},z^{*}\ \text{s.t.}\quad z^{*}$ $=Kx^{*}\quad\text{where}\quad x^{*}\in\partial f^{*}(-K^{\top}u^{*}),\,z^{*}\in\partial g^{*}(u^{*}).$
@@
$T_{1}(u)$ $=-K\partial f^{*}(-K^{\top}u)=\partial(f^{*}\circ(-K^{\top}))(u),$
@@
We need to solve the inclusion problem $0\in T_{1}(u)+T_{2}(u)$. Using DRS 9 on the primal variable $z_{k}+w_{k}$, and where the dual variable is given by $u_{k}$, the iterations may be rewritten as
@@
The resolvent of $T_{2}$ may be recognized exactly as a proximal operator $J_{\rho T_{2}}=J_{\rho\partial g^{*}}=\operatorname{prox}_{\rho g^{*}}=\operatorname{Id}-\rho\operatorname{prox}_{\rho^{-1}g}(\rho^{-1}\cdot)$. The resolvent of the first operator may also be computed as
$J_{\rho T_{1}}(v)$ $=\operatorname{prox}_{\rho f^{*}\circ(-K^{\top})}(v)$
$=v-\rho\operatorname{prox}_{\rho^{-1}[f^{*}\circ(-K^{\top})]^{*}}(\rho^{-1}v)$
$=v-\rho\operatorname*{arg\,min}_{x}\inf_{z\text{ s.t.}-Az=x}f(z)+\frac{\rho}{2}\|x-\rho^{-1}v\|^{2}$
$=v+\rho K\operatorname*{arg\,min}_{z}\left(f(z)+\frac{\rho}{2}\|Kz+\rho^{-1}v\|^{2}\right).$
@@
$\begin{cases}\hat{x}_{k+1,1}=\operatorname*{arg\,min}_{x_{1}}\left(f(x_{1})+z_{k}^{\top}(Kx_{1})+\frac{\rho}{2}\|Kx_{1}-\hat{x}_{k,2}\|^{2}\right),\\ \hat{x}_{k+1,2}=\operatorname*{arg\,min}_{x_{2}}\left(g(x_{2})-z_{k}^{\top}x_{2}+\frac{\rho}{2}\|K\hat{x}_{k+1,1}-x_{2}\|^{2}\right),\\ z_{k+1}=z_{k}+\rho(K\hat{x}_{k+1,1}-\hat{x}_{k+1,2}).\end{cases}$ (22)
@@
We observe that the minimization steps in 22 are precisely the minimizations with respect to the first two arguments of 23. Such an equivalence provides a powerful link to the theoretically simpler DRS, providing initial convergence results for ADMM with relaxation factors $\rho$ through the monotone operator framework eckstein1992douglas . In the context of imaging inverse problems, the PnP variant of the ADMM algorithm emerges by replacing the proximal operator (with respect to a nonsmooth regularizer $g$) with an off-the-shelf denoiser $D$, while minimizing the variational objective with an $\ell_{2}^{2}$ fidelity term: $\underset{x}{\min}\,\frac{1}{2}\|y-Kx\|_{2}^{2}+g(x)$. To solve this problem using ADMM, one applies the variable separation trick to reformulate the problem as
@@
$\begin{cases}x_{k+1}=\left(K^{\top}K+\rho\,\operatorname{Id}\right)^{-1}\left(K^{\top}y+\rho\,(z_{k}-u_{k})\right),\\ z_{k+1}=D\left(x_{k+1}+u_{k}\right),\\ u_{k+1}=u_{k}+\left(x_{k+1}-z_{k+1}\right).\end{cases}$ (PnP-ADMM)